Convert 8.2e-4 Molar of Mg to ppm
Answer #1

Sol:-

We know

1 mg/L = 1 ppm ( i.e part per million )

it means we have to convert Molarity ( i.e mol/L ) into mg/L  

Given

Molarity = 8.2 x 10-4 Molar = 8.2 x 10-4 mol/L i.e 8.2 x 10-4 moles of Mg present in 1 Litre of the solution.

we know that

number of moles of Mg = mass of Mg in gram / gram atomic mass of Mg

so

Mass of Mg in gram = number of moles of Mg x gram atomic mass of Mg

Mass of Mg in gram = 8.2 x 10-4 mol x 24 g/mol = 0.01968 g

Mass of Mg in milligrams = 19.68 mg  

(because 1 gram = 1000 milligrams)

Therefore

8.2 x 10-4 moles = 19.68 mg

or

8.2 x 10-4 Molar = 8.2 x 10-4 mol/L = 19.68 mg/L = 19.68 ppm ( because 1ppm = 1 mg/L)

Hence

8.2 x 10-4 Molar = 19.68 ppm